Toga flew her ultralight plane to a landing field 24 mi away. with the wind, the flight took 2/5 h. Returning against the wind, the flight took 3/4 h. Find the rate of the plane in still air and the rate of the wind.
Answer by checkley77(12569)

D=RT
24=R*.4
.4R=24
R=24/.4
R=60 MPH WITH THE WIND
24=R*.75
R=24/.75
R=32 MPH AGAINST THE WIND.
(60-32)/2=28/2=14 MPH IS THE WIND SPEED.
60-14=46 THE ULTRALIGHT SPEED.
32+14=46 DITTO.